OpenStudy (anonymous):

Find S9 for the series 1 + 5 + 25 + 125 +... i cnt be anything? bc i got 0

OpenStudy (michele_laino):

it is a geometric sequence whose first term is a1= 1 and constant q is 5

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

so: \[\Large {S_9} = {a_1}\frac{{1 - {q^9}}}{{1 - q}} = 1 \cdot \frac{{1 - {5^9}}}{{1 - 5}} = ...?\]

OpenStudy (anonymous):

488281

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Find S10 for the series 1 + 2 + 4 + 8 +... is it the same formula if its not mulitpying by 2?

OpenStudy (michele_laino):

here we have: first term= 1 constant= 2 so: \[\Large {S_{10}} = {a_1}\frac{{1 - {q^{10}}}}{{1 - q}} = 1 \cdot \frac{{1 - {2^{10}}}}{{1 - 2}} = ...?\]

OpenStudy (anonymous):

1023

OpenStudy (michele_laino):

correct!